OpenStudy (anonymous):

How many sides does a polygon with an interior angle measure of 108 degrees have? Show your work. How many sides does a polygon with an interior angle measure of 108 degrees have? Show your work. @Mathematics

OpenStudy (amistre64):

id says at least 5 but; 180 - intAng = centralAng 360/centralAng = number of sides

OpenStudy (amistre64):

in know a square has 90 and a hex has 120 and this is between them so id guess at 5

OpenStudy (anonymous):

can u show the steps please i do not understand?

OpenStudy (anonymous):

180 - 108 = 72 360/72 = 5 sides

OpenStudy (phi):

 the sum of all the turns has to be 360. the amount you turn is 180-interior angle. For a triangle, 60 degrees, so the turn is 120. For a pentagon, the interior angle is 72, and the exterior angle is 180-72= 108. Every time you turn, you turn 108 degrees and after 5 turns you get all the way around, 360.